OpenStudy (anonymous):

Determine if the trinomial is a perfect square. If so factor, if not explain why A) A fence will be built around a garden with an area of (49 + 56x + 16) . The dimensions of the garden are cx+d, where c and d are whole numbers. Find an expression for the perimeter of the garden. Find the perimeter when x=5 feet

OpenStudy (anonymous):

I'm assuming that the garden's Area = 49x^2 + 56x +16 Therefore, A = (cx +d)^2 = 49x^2 +56x +16 = (7x +4)^2 c = 7 d =4 side length s = cx +d = 39 Perimeter = 4s = 156